OpenStudy (anonymous):

(-9x^3 + 3x^2 - 15x) / (-3x)(-3x) I know it equals 6x-1 typing it in on the calculator but how do you find the quotient exactly? I have to show work but don't how the heck to even start to look for the answer

OpenStudy (anonymous):

no it is not

OpenStudy (johnweldon1993):

Yeah I'm not too sure where you are getting THAT answer....can you show us how you arrived to that?

OpenStudy (anonymous):

I'm lost and confused I typed it on this app I have on my phone call wolframalpha but I wasn't completely sure

OpenStudy (anonymous):

is it \[\frac{-9x^3 + 3x^2 - 15x}{-3x}\]?

OpenStudy (anonymous):

Yes it is

OpenStudy (johnweldon1993):

Okay so it is just one ( -3x ) on the bottom?? not ( -3x )( -3x ) as stated in the question?

OpenStudy (anonymous):

no oh I am sorry I didn't realize I made a typo. Satellite posted the correct question

OpenStudy (johnweldon1993):

Okay....so can you answer these 3 questions... what is... \[\frac{ -9x^3 }{ -3x } = ?\] \[\frac{ 3x^2 }{ -3x } = ?\] \[\frac{ -15x }{ -3x } = ?\]

OpenStudy (anonymous):

3x-x+5 right?

OpenStudy (johnweldon1993):

Very close \[3x^2 - x + 5\] Because remember....-9x^3/-3x is the same as -3x^(3 - 1) or -3x^2 But other than that...good work!

OpenStudy (anonymous):

So is that the answer and thanx

OpenStudy (johnweldon1993):

That would be the answer \[3x^2 - x + 5\]

OpenStudy (anonymous):

Thank you